A ray of light makes an angle of 10° with the horizontal above it and strikes a plane mirror which is inclined at an angle O to the horizontal. The angle O for which the reflected ray becomes vertical is
(a) 40°.         (B)50°.        (C)80°.          (D) 100°
 

Grade:12

3 Answers

Abhishek Kumar
9 years ago
Draw a ray-diagram,
Angle between the reflected and incident ray = 80 (Final ray is vertical and incidence is 10 wrt horizontal)
Angle of incidence = angle of reflected ray = 40
Angle of reflected ray with mirror = 90 – 40 = 50
So, angle with horizontal = 90 – 50 = 40
SAKSHI
5 years ago
A ray of light makes an angle of 10° with the horizontal above it and strikes a plane mirror which is inclined at an angle O to the horizontal. The angle O for which the reflected ray becomes vertical is
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your question.
Angle between the reflected and incident ray = 80 (Final ray is vertical and incidence is 10 wrt horizontal)
Angle of incidence(i) = angle of reflection(r) = 40
Angle of reflected ray made with mirror = 90 – 40 = 50
So, angle with horizontal = 90 – 50 = 400

Hope it helps.
Thanks and regards,
Kushagra